# When predict can also explain: few-shot prediction to select better neural latents

### v3

@@ 4 Good co-smoothing does not guarantee correct latents
$A_{m,l}$ $=p(z_{t+1}=l|z_{t}=m)$ $\forall\ m,l$ (6)
$B_{m,n}$ $=p(x_{n,t}=1|z_{t}=m)$ $\forall\ m,n$
$\pi_{m}$ $=p(z_{0}=m)$ $\forall\ m$
@@
i.e., smoothing, computed exactly with the forward-backward algorithm (Barber, 2012). Note that although $z$ is the latent state of the HMM, we use its posterior probability mass function $\bm{\xi}_{t}$ as the relevant intermediate representation. To make predictions of the rates of held-out neurons for co-smoothing we compute:
@@
As a teacher, we constructed a 4-state model of a noisy chain $A_{m,l}\propto\mathbb{I}[l=(m+1)\mod M]+\epsilon$, with $\epsilon=1e-2$, $\pi=\frac{1}{M}$, and $B_{m,n}\sim\text{Unif}(0,1)$ sampled once and frozen (Fig. 2, left). We generated a dataset of observations from this teacher (see appendix I). We trained $400$ students with $4-15$ states on the same teacher data using gradient-based methods (see appendix B). All students had high co-smoothing scores, with some variance, and a trend for large students to perform better. Consistent with our hypothesis, the ability to decode the teacher from the student varied little, and was highly correlated to the co-smoothing score (Fig. 1D). In contrast, the ability to decode the student from the teacher displayed a large variability, and little correlation to the co-smoothing score (Fig. 1E). See appendix C for details of the regression.
What is it about a student model, that produces good co-smoothing with the wrong latents? We consider the HMM transition matrix for the teacher and two exemplar students – named “Good” and “Bad” (marked by green and red arrows in Fig. 1CD) – and visualise their states and transitions using graphs in Fig. 2. The teacher is a cycle of 4 steps. The good student also contains such a cycle (orange), and the initial distribution $\pi$ is restricted to that cycle, rendering the other states irrelevant. In contrast, the bad student also contains this cycle (orange), but the $\pi$ distribution is not consistent with the cycle leading to an extraneous branch converging to the cycle, as well as a departure from the main cycle (both components in dark colour). Note that this does not interfere with co-smoothing, because the emission probabilities $B_{m,n}$ of the extra states are consistent with true states, i.e., the emission matrix conceals the extraneous dynamics.
Figure 2: Visualisations of HMMs: the ground truth T along with two representative extreme student models. Circles represent states, and arrows represent transitions. Circle area and edge widths, reflect fraction of visitations or volume of traffic after sampling the HMM over several trials, each trial, a sequence of $T$ time steps. The colours also reflect the same quantity – brighter for higher traffic. Edges with values below $0.02$ are removed for clarity. The teacher ($M=4$) is a $4$-cycle. Note the prominent 4-cycles (orange) present in the good student ($M=10$), and the bad student ($M=8$). In the good student, the extra states are seldom visited, whereas in the bad student there is significant extraneous dynamics involving these states (dark arrows).
5 Few-shot prediction selects better models
Because our objective is to obtain latent models that are close to the ground truth, the co-smoothing prediction scores described above are not satisfactory. Can we devise a new prediction score that will be correlated with ground truth similarity? The advantage of prediction benchmarks is that they can be optimized, and serve as a common language for the community as a whole to produce better algorithms (Deng et al., 2009).
We suggest few-shot co-smoothing as a complementary prediction score to co-smoothing, to be used on models with good scores on the latter. Similarly to standard co-smoothing, the functions $g$ and $f$ are trained using all trials of the training data (Fig. 3A). The key difference is that a separate group of $N^{\text{$k$-out}}$ neurons is set aside, and only $k$ trials of these neurons are used to estimate a mapping $g^{\prime}:{\bm{Z}}_{t,:}\mapsto{\bm{R}}_{t,\text{$k$-out}}$ (Fig. 3B), similar to $g$ in equation 2. The neural LVM $(f,g,g^{\prime})$ is then evaluated on both the standard co-smoothing $\mathcal{Q}$ using $g\circ f$ and the few-shot version $\mathcal{Q}^{k}$ using $g^{\prime}\circ f$ (Fig. 3C).
Figure 3: Co-smoothing and few-shot co-smoothing; a composite evaluation framework for Neural LVMs. A. The encoder $f$ and decoder $g$ are trained jointly using held-in and held-out neurons. B. A separate decoder $g^{\prime}$ is trained to readout $k$-out neurons using only $k$ trials. Meanwhile, $f$ and $g$ are frozen. C. The neural LVM is evaluated on the test set resulting in two scores: co-smoothing $\mathcal{Q}$ and $k$-shot co-smoothing $\mathcal{Q}^{k}$.
This procedure may be repeated several ($s$) times independently on resampled sets of $k$ trials, giving $s$ estimates of $g^{\prime}$, each yielding a score $\mathcal{Q}^{k}$ for each $k$-set. For small $k$, the $\mathcal{Q}^{k}$s tend to be highly variable. Thus we compute and report the average score $\langle\mathcal{Q}^{k}_{\text{S}}\rangle$ over the $s$ resamples for each student S. Practical advice on how to choose the value of $k$ and $s$ is given in appendix H.
To show the utility of the newly-proposed prediction score, we return to the same HMM students from Fig. 1. For each student, we evaluate $\langle\mathcal{Q}^{k}_{\text{S}}\rangle$. This involves estimating the bernoulli emission parameters $\hat{\bm{B}}_{m,\text{$k$-out}}$, given the latents $\xi_{t,m}^{(i)}$ using equation 44 and then generating rate predictions for the $k$-out neurons using equation 8. First, we see that it provides new information on the models, as it is not correlated with standard co-smoothing (Fig. 4A). We also show that it is not simply a harder version of co-smoothing (appendix D). We are only interested in models that have good co-smoothing, and thus select students satisfying $\mathcal{Q}_{\text{S}}>\mathcal{Q}_{\text{T}}-\epsilon$, choosing $\epsilon=10^{-3}$. For these students, we see that despite having very similar co-smoothing scores, their $k$-shot scores $\langle\mathcal{Q}^{k}_{\text{S}}\rangle$ are highly correlated with the ground truth measure $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ (Fig. 4B). Taken together, these results suggest that the combined objective of maximising $\mathcal{Q}_{\text{S}}$ and $\langle\mathcal{Q}^{k}_{\text{S}}\rangle$ simultaneously – both prediction based objectives – yields models achieving low $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ and $\mathcal{D}_{\text{T}\rightarrow\text{S}}$, a more complete notion of model similarity to the ground truth.
Figure 4: Few-shot prediction selects better models. A. Student models with high co-smoothing have highly variable $6$-shot co-smoothing and uncorrelated to co-smoothing. B. For the set of students with high co-smoothing, i.e., satisfying $\mathcal{Q}_{\text{S}}>\mathcal{Q}_{\text{T}}-10^{-3}$, $6$-shot co-smoothing to held-out neurons is negatively correlated with decoding error from teacher-to-student. Following Fig. 4C,D dashed lines represent the ground truth, green and red arrows represent “Good” and “Bad” models (Fig. 2).
6 Why does few-shot work?
The example HMM students of Fig. 2 can help us understand why few-shot prediction identifies good models. The students differ in that the bad student has more than one state corresponding to the same teacher state. Because these states provide the same output, this feature does not hurt co-smoothing. In the few-shot setting, however, the output of all states needs to be estimated using a limited amount of data. Thus the information from the same amount of observations has to be distributed across more states. We make this data efficiency argument more precise in three settings: HMMs, linear regression and prototype learning.
Consider a student-teacher scenario as in section 4. We let $T=2$ and use a stationary teacher $z^{(i)}_{1}=z^{(i)}_{2}$. Now consider two examples of inferred students. To ensure a fair comparison, we use two latent states for both students. In the good student, $\xi$, these two states statistically do not depend on time, and therefore it does not have extraneous dynamics. In contrast, the bad student, $\mu$, uses one state for the first time step, and the other for the second time step. A particular example of such students is:
$\xi_{t}=\begin{bmatrix}0.5&0.5\end{bmatrix}^{T}\ t\in\{1,2\}$ (9)
We can now evaluate the maximum likelihood estimator of the emission matrix $\hat{{\bm{B}}}$ from $k$ trials for both students, and compare it to the teacher ${\bm{B}}^{*}$. In the case of bernoulli HMMs the maximum likelihood estimate of $g^{\prime}$ given a fixed $f$ and $k$ trials has a closed form. For a single output neuron, it is:
$\begin{array}[]{ccc}\hat{B}(\xi)=\frac{0.5(C_{1}+C_{2})}{0.5kT}&&\hat{B}_{1}(\mu)=\frac{C_{1}}{k}\\ \hat{B}_{2}(\xi)=\frac{0.5(C_{1}+C_{2})}{0.5kT}&&\hat{B}_{2}(\mu)=\frac{C_{2}}{k}\\ \end{array}$ (12)
where $C_{t}$ is the number of times $x=1$ at time $t$ in $k$ trials. We see that $C_{t}$ is a sum of $k$ i.i.d. Bernoulli random variables (RVs) with the teacher parameter $B^{*}$, for both $t=1,2$. We see that for both students $\hat{{\bm{B}}}$ are unbiased, but the bad student’s $\hat{{\bm{B}}}$ has higher variance.
Using a Taylor expansion of the loss around the teacher emission matrix (appendix N), we obtain:
$\mathbb{E}_{\hat{B}_{k}}L_{\xi}\left(\hat{B}_{k}\left(\xi\right)\right)$ $\approx L(B^{*})-\frac{1}{2k},$ (13)
$\mathbb{E}_{\hat{B}_{k}}L_{\mu}\left(\hat{B}_{k}\left(\mu\right)\right)$ $\approx L(B^{*})-\frac{1}{k}.$ (14)
Thus, the good teacher $\xi$ has a better performance for low $k$.
We can construct a similar scenario for linear regression, with continuous latents (appendix L). The latents have a signal dimension and several noise dimensions, corresponding to the extraneous dynamics. In this case, the variance does not discriminate between students, but the bias does – still leading to higher error for extraneous dynamics in low $k$. Crucially, if $k$ larger than a critical value, the effect vanishes.
Finally, we use prototype learning for binary classification on continuous latents (appendix M). Here too, we show that variance underlies the ability of low $k$ to discriminate models. In this setting, however, the effect gradually diminishes with $k$.
7 SOTA LVMs on neural data
In section 4 we showed that models with near perfect co-smoothing may possess latents with extraneous dynamics. We established this in a synthetic student-teacher setting with simple HMM models.
To show the applicability in more realistic scenarios, we trained several models from two SOTA architectures, LFADS (Sedler & Pandarinath, 2023; Pandarinath et al., 2018; Keshtkaran et al., 2022), a variational autoencoder (Kingma, 2013), and STNDT (Le & Shlizerman, 2022; Ye & Pandarinath, 2021), a transformer (Nguyen & Salazar, 2019; Huang et al., 2020), on mc_maze_20 consisting of neural activity recorded from monkeys performing a maze solving task (Churchland et al., 2010), curated by (Pei et al., 2021). The 20 indicates that spikes were binned into $20ms$ time bins. We evaluate co-smoothing on a test set of trials and define the set of models with the best co-smoothing (appendix F and I).
An integral part of LFADS and STNDT training is the random hyperparameter sweep which generates several candidate solutions to the optimization problem equation 5.
With each model $f_{u}$, we infer latents evaluated over a fixed set of test trials $\mathcal{X}^{\text{test}}$, using equation 1.
In the HMM case, we had ground truth that enabled us to directly compare the student latent to that of the teacher. With real neural data we do not have this privilege. To nevertheless reveal the presence or absence of extraneous dynamics, we instead compare the models to each other. The key idea is that all models contain the teacher latent, because they have good co-smoothing. One can then imagine that each student contains a selection of several extraneous features. The best student is the one containing the least such features, which would imply that all other students can decode its latents, while it cannot decode theirs. We therefore use cross-decoding among student models as a proxy to the ground truth.
Instead of computing $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ and $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ as in section 4 we perform cross-decoding from latents of model $u$ to model $v$ ($\mathcal{D}_{u\rightarrow v}$) for every pair of models $u$ and $v$ using linear regression and evaluating an $R^{2}$ score for each mapping (appendix F). In Fig. 19 the results are visualised by a $U\times U$ matrix with entries $\mathcal{D}_{u\rightarrow v}$ for all pairs of models $u$ and $v$.
We hypothesize that the latents $z_{u}$ contain the information necessary to output good rate predictions $r$ that match the outputs plus the arbitrary extraneous dynamics. This former component must be shared across all models with high $\mathcal{Q}$, whereas the latter could be unique in each model – or less likely to be consistent in the population. The ideal model $v^{*}$ would have no extraneous dynamics therefore, all the other models should be able to decode to it with no error, i.e., $\mathcal{D}_{u\rightarrow v^{*}}=0\ \forall\ u$. Provided a large and diverse population of models only the ‘pure’ ground truth would satisfy this condition. To evaluate how close is a model $v$ to the ideal $v^{*}$ we propose a simple metric: the column average $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$. This will serve as proxy for the distance to ground truth, analogous to $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ in Fig. 4. We validate this procedure using the student-teacher HMMs in appendix G, where we show it is highly correlated to ground truth, and as correlated to few-shot as the SOTA models.
Figure 5: Cross-decoding as a proxy for distance to the ground truth in near-SOTA models. 120 STNDT models were trained on the mc_maze_20 dataset then selected for high-cosmoothing (appendix F). Top: The latents of each pair of models were decoded from one another, and the decoding error is shown in the matrices. Good models are expected to be decoded from all other models, and hence have low values in their corresponding columns. Bottom: Trajectories of two STNDT models, with the lowest (left and in the green box, best model$:=\text{argmin}_{v}\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$) and highest (right and in the red box, the worst model$:=\text{argmax}_{v}\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$) column averaged cross-decoding errors, projected onto their leading two principal components. Scores for these models are indicated in Fig. 6 by the arrows. Each trace is the trajectory for a single trial, starting at a green dot and ending at a red dot. The same analysis was performed on LFADS models (appendix O).
Figure 6: Few-shot scores correlate with the proxy of distance to the ground truth. Several models of two architectures (LFADS top, STNDT bottom) were trained on neural recordings from monkeys performing a maze task, the mc_maze_20 benchmark (Churchland et al., 2010; Pei et al., 2021). Distance to ground truth was approximated by the cross-decoding column average $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$ (Fig. 19). Few-shot ($k=128$) co-smoothing scores (left) negatively correlate with $\mu$, while regular co-smoothing (right) does not. Green and red arrows indicate the extreme models whose latents are visualised in Fig. 19 matched by box/arrow colours. $\mathcal{Q}_{v}$ values may be compared against an EvalAI leaderboard (Pei et al., 2021). Note that we evaluate using an offline train-test split, not the true test set used for the leaderboard scores, for which held-out neuron data is not directly accessible.
Having developed a proxy for the ground truth we can now correlate it with the few-shot regression to held-out neurons. Fig. 6 shows a negative correlation for both architectures, similar to the HMM examples above. As an illustration of the latents of different models, Fig. 19 shows the PCA projection of several trials from two different models. Both have high co-smoothing scores (LFADS: $0.3647$,$0.3643$, STNDT: $0.3488,0.3495$), but differ in their cross-decoding column average $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$. Note the somewhat smoother trajectories in the model with higher few-shot score. It is also possible to cross-decode across the two populations, as shown in Appendix K.
8 Discussion
Latent variable models aim to infer the underlying latents using observations of a target system. We showed that co-smoothing, a common prediction measure of the goodness of such models cannot discriminate between certain classes of latents. In particular, extraneous dynamics can be invisible to such a measure.
We suggest a complementary prediction measure: few-shot co-smoothing. Instead of directly regressing from held-in to held-out neurons as is done to evaluate co-smoothing, we distinguish the encoder and the decoder. To evaluate the trained model we substitute the decoder with a new decoder estimated using only a ‘few’ ($k$) number of trials. The rate predictions provided by the few-shot decoder are evaluated the same as in standard co-smoothing. Using synthetic datasets and HMM models, we show numerically and analytically that this measure correlates with the distance of model latents to the ground truth.
We demonstrate the applicability of this measure to real world neural datasets, with SOTA architectures. This required developing a new proxy to ground truth – cross decoding. For each pair of SOTA models that we obtained, we performed a linear regression across model latents, provided identical input data. Models with extraneous dynamics showed up as a bad target latent on average, and vice versa. Finally we show that these two characterisations of extraneous dynamics are correlated. An interesting extension would be to use this new metric as another method to select good models. The computational cost is high, because it requires training a population of models and comparing between all of them. It is also less universal and standardised than few-shot co-smoothing, as it is dependent on a specific ‘jury’ of models. The HMM results, however, show that it is more correlated to ground truth than the few-shot.
While the combination of student-teacher and SOTA results put forth a compelling arguement, we address here a few limitations of our work. Firstly, our SOTA results use only one of the datasets in the benchmark suite (Pei et al., 2021). With regard to the few-shot regression, while the bernoulli HMM scenario has a closed form solution: the maximum likelihood estimate, the poisson GLM regression for the SOTA models is optimised iteratively and is sensitive to the l2 hyperparameter alpha. In our results we select $k$ and $\alpha$ that distinguish models in our candidate model sets giving moderate/high few-shot scores for some models and low scores to others. This is an empirical choice that must be made for each dataset and model-set. The few-shot training of $g^{\prime}$ is computationally inexpensive and may be thus can evaluated over a range of values to find the ideal ones.
Overall, our work advances latent dynamics inference in general and prediction frameworks in particular. By exposing a failure mode of standard prediction metrics, we can guide the design of inference algorithms that take this into account. Furthermore, the few-shot prediction can be incorporated into existing benchmarks and help guide the community to build models that are closer to the desired goal of uncovering latent dynamics in the brain.
Appendix A Additional discussion
Why use HMMs? The SOTA LVMs are typically deep-neural networks. HMMs are much too simple.
Real neural data is trajectories across time, and therefore a teacher model should be a dynamical system. HMMs were chosen as a tractable dynamical system to show our claims. We highlight specific properties of HMMs that are relevant to our use:
  • HMMs are perhaps the simplest stochastic dynamical system, making analysis and visualisation possible, e.g., the graphs shown in Fig. 2 show that the students contain the teacher’s dynamics – a valuable insight,
  • HMMs are expressive enough to capture real neural dynamics. Indeed, HMMs appear on the leaderboard for the mc_maze_20 benchmark (see here). Although not the leading models, they outperform other baselines.
  • They may be used for both data generation and inference of latents – perhaps the simplest Latent Variable Models for time-series data. By constraining the student and teacher to the same class of models, we are justified in directly comparing the two.
How to choose $k$?
As $k$ is increased, the ability of the few-shot metric to discriminate extraneous dynamics decreases. The exact manner depends on the setting (see theoretical analysis in appendices L, M, N). On the other hand, decreasing $k$ too much increases noise, and eventually precludes reaching any conclusion. Appendix H provides practical advice on the choice of $k$ in real-world scenarios.
Why use cross-decoding as a proxy, and not directly to evaluate models instead of few-shot co-smoothing?
In the HMM setting, we see that indeed cross-decoding is more correlated to ground truth than few-shot co-smoothing (figures 4 and 10). Nevertheless, it is qualitatively different from few-shot co-smoothing and has several disadvantages. First, the computational cost is high because it requires training a population of models and comparing between all directed pairs of them. Furthermore, it is also less universal and standardised than few-shot co-smoothing, as it is dependent on a specific ‘jury’ of models – the population itself. On the other hand, few-shot co-smoothing is computed for each model independently. Furthermore, its close relation to co-smoothing, the current standard for benchmarking neural LVMs, makes it easier to incorporate into existing benchmarking pipelines. For these reasons, we identify few-shot co-smoothing as the better option for evaluation.
Scope of applicability. How general are the conclusions from the HMM student-teacher experiments? There is a large gap between HMMs and modern SOTA models discussed in the main text.
Both our theoretical and experimental results extend beyond HMMs. Theoretically, we show that few-shot regression is sensitive to extraneous dynamics in three different settings (appendices L, M, N). In all three settings, the true and extraneous latent representations are distinguished by poorer generalisation performance of the few-shot regression for the extraneous model at some finite $k$ and the difference vanishes as $k\to\infty$. Experimentally, we replicate the HMM results in a continuous setting – Linear Gaussian State Space Models (LGSSMs) in appendix E. The SOTA results in the main text are an additional experimental validation.
If you want simple students, why not use regularization on the latents?
Previous work also noted that students can predict well despite having the wrong latents (Sedler et al., 2022). To alleviate this, regularization on the latents was suggested (Sedler et al., 2022; Versteeg et al., 2023). Our approach is different in that we remain within the latent-agnostic prediction framework. Despite not requiring access to the latents, few-shot prediction encourages models without extraneous dynamics. This implicit form of regularization has the advantage of being easily integrated within benchmarks and allowing optimization of algorithms without restriction.
Does few-shot co-smoothing replace regular co-smoothing?
No. We propose few-shot co-smoothing as a secondary metric to accompany regular co-smoothing. The latter encourages the student to contain as much of the teacher dynamics as possible, while the former encourages the student to have as little extraneous dynamics differing from that of the teacher. We show that student models with good performance in both metrics, are the students most similar to the teacher.
Appendix B Hidden Markov Model training
@@
HMMs are traditionally trained with expectation maximisation, but they can also be trained using gradient-based methods. We focus here on the latter as these are used ubiquitously and apply to a wide range of architectures. We use an existing implementation of HMMs with differentiable parameters: dynamax – a library of differentiable state-space models built with jax.
@@
To find the minimum we do full-batch gradient descent on $L$, using dynamax together with the Adam optimiser (Kingma & Ba, 2014) .
Appendix C Decoding across HMM latents: fitting and evaluation
@@
In section 4 and Fig. 1, the data $X$ is sampled from a single teacher HMM, T, and we evaluate $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ and $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ for each student notated simply as S.
Appendix D Few-shot co-smoothing is not simply hard co-smoothing
The few-shot benchmark is a more difficult one than standard co-smoothing. Thus, it might seem that any increase in the difficulty of the benchmark will yield similar results. To show this is not the case, we use standard co-smoothing with fewer held-in neurons (Fig. 7). The score is lower (because it’s more difficult), but does not discriminate models.
Figure 7: Making co-smoothing harder does not discriminate between models. Top three: Increasing the number of held out neurons from $N^{\text{out}}=50$ to $N^{\text{out}}=100$. First two panels: Same as main text Fig. 1CD. Lower panel: Same as main text Fig. 4B. Bottom three: Decreasing the number of held-in and held-out neurons to $N^{\text{in}}=5$, $N^{\text{out}}=5$, $N^{\text{$k$-out}}=50$. Panels as in top row. The score does decrease because the problem is harder, but co-smoothing is still not indicative of good models while few-shot is.
Appendix E Student-teacher results in Linear Gaussian State Space Models
We demonstrate that our results are not unique to the HMM setting by simulating another simple scenario: linear gaussian state space models (LGSSM), i.e., Kalman Smoothing.
The model is defined by by parameters $({\bm{\mu}}_{0},{\bm{\Sigma}}_{0},{\bm{F}},{\bm{G}},{\bm{H}},{\bm{R}})$. A major difference to HMMs is that the latent states $\bm{z}\in\mathbb{R}^{M}$ are continuous. They follow the dynamics given by:
${\bm{z}}_{0}$ $\sim\mathcal{N}({\bm{\mu}}_{0},{\bm{\Sigma}}_{0})$ (20)
${\bm{z}}_{t}$ $\sim\mathcal{N}({\bm{F}}{\bm{z}}_{t-1}+{\bm{b}},{\bm{G}})$ (21)
${\bm{x}}_{t}$ $\sim\mathcal{N}({\bm{H}}{\bm{z}}_{t}+{\bm{c}},{\bm{R}})$ (22)
As with HMMs we use a teacher LGSSM with $M=4$, with parameters chosen randomly (using the dynamax defaults) and then fixed. Student LGSSMs are also initialised randomly and optimised with Adam (Kingma & Ba, 2014) to minimise negative loglikelihood on the training data (see table 1 for dimensions of the synthetic data set). $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ and $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ is computed with linear regression (sklearn.linear_model.LinearRegression) and predictions are evaluated against the target using $R^{2}$ (sklearn.metrics.r2_score). We define $\mathcal{D}_{u\rightarrow v}:=1-(R^{2})_{u\rightarrow v}$. Few-shot regression from $\bm{z}$ to $\bm{x}^{\text{$k$-out}}$ is also performed using linear regression.
In line with our results with HMMs (sections 4 and 5), in Fig. 8 we show that that among the models with high test loglikelihood ($>-55$), $\mathcal{D}_{\text{S}\rightarrow\text{T}}$, but not $\mathcal{D}_{\text{T}\rightarrow\text{S}}$, is highly correlated to test loglikelihood, while $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ shows a close relationship to Average 10 shot MSE error.
Figure 8: Left to right: Student-teacher results for Linear Gaussian State Space Models. We report loglikelihood instead of co-smoothing, and $k$-shot MSE instead of $k$-shot co-smoothing.
Appendix F Analysis of SOTA models
@@
We denote the set of high co-smoothing models as those satisfying $\mathcal{Q}_{\text{model}}>\mathcal{Q}_{\text{best model}}-\epsilon$, choosing $\epsilon=5\times 10^{-3}$ for LFADS and $\epsilon=1.3\times 10^{-2}$ for STNDT. :=$\mathcal{F}:=\{(f_{u},g_{u})\}_{u=1}^{U}$, the encoders and decoders respectively. Note that both architectures are deep neural networks given by the composition $g\circ f$, and the choice of intermediate layer whose activity is deemed the ‘latent’ ${\bm{Z}}$ is arbitrary. Here we consider $g$ the last ’read-out’ layer and $f$ to represent all the layers up-to $g$. $g$ takes the form of Poisson Generalised Linear Model (GLM), a natural and simple choice for the few-shot version $g^{\prime}$. To this end, we use sklearn.linear_model.PoissonRegressor. The poisson regressor has a hyperparameter alpha, the amount of l2 regularisation. For the results in the main text, $\langle\mathcal{Q}_{v}^{k}\rangle$ in Fig. 6, we select $\alpha=10^{-3}$.
To perform few-shot co-smoothing, we partition the train data into several subsets of $k$ trials. To implement this in a standarised way, we build upon the nlb_tools library (appendix J). This way we ensure that all models are trained and tested on identitical partitions.

### v4

@@ Methods
$A_{m,l}$ $=p(z_{t+1}=l|z_{t}=m)\$$\forall\ m,l$ (14)
$B_{m,n}$ $=p(x_{n,t}=1|z_{t}=m)\$$\forall\ m,n$
$\pi_{m}$ $=p(z_{0}=m)\$$\forall\ m$
@@
i.e., smoothing, computed exactly with the forward-backward algorithm [51]. Note that although $z$ is the latent state of the HMM, we use its posterior probability mass function $\bm{\xi}_{t}$ as the relevant intermediate representation because it reflects a richer representation of the knowledge about the latent state than a single discrete state estimate. To make predictions of the rates of held-out neurons for co-smoothing we compute:
@@
As a teacher, we constructed a 4-state model of a noisy chain $A_{m,l}\propto\mathbb{I}[l=(m+1)\mod M]+\epsilon$, with $\epsilon=1e-2$, $\pi=\frac{1}{M}$, and $B_{m,n}\sim\text{Unif}(0,1)$ sampled once and frozen (Fig. 2, left). We generated a dataset of observations from this teacher (see Table 1).
For each student, we evaluate $\langle\mathcal{Q}^{k}_{\text{S}}\rangle$. This involves estimating the bernoulli emission parameters $\hat{\bm{B}}_{m,\text{$k$-out}}$, given the latents $\xi_{t,m}^{(i)}$ using equation 27 and then generating rate predictions for the $k$-out neurons using equation 16.
HMM training
@@
HMMs are traditionally trained with expectation maximisation, but they can also be trained using gradient-based methods. We focus here on the latter as these are used ubiquitously and apply to a wide range of architectures. We use an existing implementation of HMMs with differentiable parameters: dynamax [52] – a library of differentiable state-space models built with jax.
@@
To find the minimum we do full-batch gradient descent on $L$, using dynamax together with the Adam optimiser [50] .
Decoding across HMM latents
@@
Since HMM latents are probability mass functions, we do not do use linear regression to learn the mappings across model latents. Instead we perform a multinomial regression from $(\bm{\xi}^{(i)}_{t})_{u}$ to $(\bm{\xi}^{(i)}_{t})_{v}$.
@@
where $W\in\mathbb{R}^{M(v)\times M(u)}$, $\bm{b}\in\mathbb{R}^{M(v)}$ and $\sigma$ is the softmax. During training we sample states from the target PMFs $(z_{t}^{(i)})_{v}\sim(\bm{\xi}^{(i)}_{t})_{v}$ thus arriving at a more well-known problem scenario: classification of $M(v)$-classes. We optimize $W$ and $\bm{b}$ to minimise a cross-entropy loss to the target $(\hat{z}_{t}^{(i)})_{v}$ using the fit() method of sklearn.linear_model.LogisticRegression.
@@
In section Good co-smoothing does not guarantee correct latents and Fig. 1, the data $X$ is sampled from a single teacher HMM, T, and we evaluate $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ and $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ for each student notated simply as S.
Analysis of LVMs without access to ground truth
@@
We denote the set of high co-smoothing models as those satisfying $\mathcal{Q}>0.034$ for Fig. 4 and $\mathcal{Q}>0.8\times\mathcal{Q}_{\text{best model}}$ in Fig. 7, $\mathcal{F}:=\{(f_{u},g_{u})\}_{u=1}^{U}$, the the encoders and decoders respectively. Note that STNDT is a deep neural network given by the composition $g\circ f$, and the choice of intermediate layer whose activity is deemed the ‘latent’ $\bm{Z}$} is arbitrary. Here we consider $g$ the last ’read-out’ layer and $f$ to represent all the layers up-to $g$.
Few-shot co-smoothing
To perform few-shot co-smoothing, we learn $g^{\prime}$, which takes the same form as $g$, a Poisson Generalised Linear Model (GLM) for each held-out neuron. We use sklearn.linear_model.PoissonRegressor, which has a hyperparameter alpha, the amount of l2 regularisation. For the results in the main text, $\langle\mathcal{Q}^{k\text{-shot}}\rangle$ in Fig. 7, we select $\alpha=10^{-3}$. We partition the training data into several random subsets of $k$ trials and train an independently initialised GLM on each subset. Each GLM is then evaluated on a fixed test set of trials (Fig. 3), yielding a score for each subset. We report the mean over $\lfloor 5\times S^{\text{train}}/k\rfloor$ such repetitions, $\langle\mathcal{Q}^{k\text{-shot}}\rangle$, along with the standard error of the mean (error bars in Fig. 4, Fig. 7). Scores are more variable at small $k$, so we need more repetitions to better estimate the average score. To implement this in a standarised way, we incorporate this chunking of data into several subsets in the nlb_tools library (S2 Appendix.). This way we ensure that all models are trained and tested on identitical subsets. We report the compute-time for few-shot co-smoothing in S3 Appendix..
Cross-decoding